OpenStudy (anonymous):

Estimate by first rounding each number to the nearest 1/2 -4 1/6*(-1 14/15 A. -12 B. -8 C. 8 D. 12

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

-4 1/6*(-1 14/15) -25/6*(-29/15)=5/6*29/3=145/18 = 8.1

OpenStudy (mathmate):

Estimate by \(first\ rounding\) each number to the nearest 1/2 -4 1/6 => -4 -1 14/15 => -2 What is -4\(\times\)(-2) ?

OpenStudy (anonymous):

is the answer is 8

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

That is correct. Remember that 8 is an estimate of the product, not the exact answer, because we approximated the multiplicands to the nearest half as instructed.

OpenStudy (anonymous):

Thank you. Could you plz help me with a few more.

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

You're welcome! Unfortunately I have to attend to something else in the next few hours. Post your questions, and if they are not answered, I will pick them up when I return.

OpenStudy (anonymous):

i only have 5 question all together and i am on 3

OpenStudy (anonymous):

Amit wants to estimate 1 1/9+3 6/7 . To estimate, he simplifies the expression 1+4 . Is Amit’s estimate a close estimate, an underestimate, or an overestimate? A. underestimate B. overestimate C. close estimate

OpenStudy (anonymous):

@party_girl @iGreen. @iGreen

OpenStudy (anonymous):

@KatnissMallark

OpenStudy (anonymous):

i think close estimate

OpenStudy (anonymous):

@BethLM39

OpenStudy (anonymous):

I'd agree

OpenStudy (anonymous):

A trail is 3 2/10 km long. Cassidy has run 9/10 of the distance of the trail. About how far has Cassidy run? Estimate by first rounding each number to the nearest 1/2 . A. 2km B. 3km C. 4km D. 7km

OpenStudy (anonymous):

Round 3 2/10 and 9/10 to the nearest half

OpenStudy (anonymous):

I do not know how

OpenStudy (anonymous):

@KatnissMallark @iGreen.

OpenStudy (anonymous):

3 or 3 1/2. Which is closer to 3 2/10?

OpenStudy (anonymous):

i think a

OpenStudy (mathmate):

Rounding to half is done as follows: 1. multiply the number by two. 2. round to the nearest integer. 3. Divide the result by two. Example: round 5 1/3 to the nearest half 1. 5 1/3 *2 = 10 2/3 2. nearest integer is 11 3. divide by two to get 5 1/2

OpenStudy (anonymous):

a?

OpenStudy (mathmate):

A trail is 3 2/10 km long. Cassidy has run 9/10 of the distance of the trail. Round 3 2/10 to nearest half: multiply by 2=6 4/10, round to nearest integer = 6 now divide by two to give 3. For 9/10, round to nearest half: multiply by 2 = 18/10 round to nearest integer to give 2, then divide by two to give 1. Now multiply the two results!